OpenStudy (anonymous):

Given b = 12, c = 15 and A = 60° in triangle ABC, use the Law of Cosines to solve for a. Fill in the blank(s) to complete each step. Round your answer to the nearest hundredth. a2 = --- + 152 - 2 --- (15) cos ---° a2 = 144 + 225 - 360(---) a2 = --- a = ---

OpenStudy (anonymous):

@zachleestone

OpenStudy (anonymous):

i have to fill the --- up

OpenStudy (perl):



OpenStudy (perl):

You can use the equation b^2 + c^2 - 2b*c*cos(60) = a^2

OpenStudy (anonymous):

a = 16.82 right? @perl

OpenStudy (perl):

not quite

OpenStudy (anonymous):

im just bad at these stuff:(

OpenStudy (anonymous):

357?

OpenStudy (perl):

that was a typo

OpenStudy (anonymous):

u mean?

OpenStudy (perl):

I left out 15 , but i fixed it

OpenStudy (anonymous):

oh what about a?

OpenStudy (perl):

If we substitute b = 12 c = 15 theta = 60 degrees into the equation b^2 + c^2 - 2ab cos (theta) = a^2 12^2 + 15^2 - 2*12*15cos(60 degrees) = a^2 144 + 225 - 2*12*15*1/2 = a^2 144 + 225 - 360*cos(60) = a^2 144 + 225 - 360 *1/2 = a^2 369 - 180 = a^2 189 = a^2

OpenStudy (anonymous):

oh ok thnx

OpenStudy (perl):

a = sqrt(189) = sqrt( 9 * 21) = sqrt(9)*sqrt(21) = 3 sqrt(21)